OpenStudy (anonymous):

Find the slope-intercept for the line with the indicated slope and y-intercept. slope -4 ; (0,-3/5)

OpenStudy (ajprincess):

Let the equation of the line be y=mx+c Here m-slope c-y_intercept m is known and to find c plug in the value y=-3/5, x=0 and m=-4 in y=mx+c. Then to find the equation plug in the value of c u get and m=-4 in y=mx+c. Can u do it @channie2003?

OpenStudy (anonymous):

i tried is it that c= -3/5

OpenStudy (ajprincess):

yup u r right:) Nw can u find ur eqn?

OpenStudy (anonymous):

-3/5=-4*0+-3/5

OpenStudy (ajprincess):

hmm no. it is y=-4x-(3/5) getting it?

OpenStudy (anonymous):

ok yes now i get it thank you

OpenStudy (ajprincess):

yw:)